HERE IS ANOTHER ONE I AM GOING CRAZY
THE
+ SAD
---------
BOOK A. THE NUMBER IS 0 APPEARS THE MOST.
B. E< 5
C. D> 5
D.THE IS A PERFECT SQUARE.
E. E + D < 10 HOW DO YOU DO THIS?

Answer by Edwin McCravy(3591)


HERE IS ANOTHER ONE I AM GOING CRAZY
   THE
 + SAD
---------
BOOK        (A). THE NUMBER IS 0 APPEARS THE MOST.
            (B). E< 5
            (C). D> 5 
            (D). THE IS A PERFECT SQUARE.
            (E). E + D < 10                      HOW DO YOU DO THIS?


Something is wrong. If THE is a perfect square, as (D) requires, then
these are the only 5 possibilities. 

       THE + SAD = BOOK
#1.    324 + 685 = 1009    
#2.    576 + 432 = 1008
#3.    625 + 379 = 1004
#4.    625 + 384 = 1009

However 
#1 violates (C)
#2 violates (B) and (C)
#3 violates (E)
$4 violates (B) and (C)

If you'll change requirement (C) to D > 6 instead of D > 5,
then #1 would be the solution.
Otherwise there is no solution.
Sorry.

Edwin